OpenStudy (anonymous):

find the product of (x+3)(x-3)

OpenStudy (anonymous):

Firsts Inners Outers Lasts. X•x, -3•x, 5•x, -3•5= x^2 - 3x + 5x - 15. Which is x^2 +2x - 15

OpenStudy (anonymous):

:) ok

HanAkoSolo (jamierox4ev3r):

what you do here, as damara tried to explain, is FOIL (x+3) (x-3) F irst x*x=x^2 O utside x*-3=-3x I nside 3*x=3x L ast 3*-3=-9 Therefore, your final equation is x^2-9

HanAkoSolo (jamierox4ev3r):

Does that make sense? @bangbang559 ?

OpenStudy (ranga):

You can also make use of the identity: (A+B)(A-B) = A^2 - B^2

OpenStudy (anonymous):

thank you so much

HanAkoSolo (jamierox4ev3r):

np :)